Write the equation of the quadratic whose one root is 2 - 5i
Answer by scott8148(6628)

complex roots occur in conjugate pairs, so the other root is 2 + 5i

if r is a root, then x-r is a factor

(x-2-5i)(x-2+5i) = x^2-2x+5ix-2x+4-10i-5ix+10i-25i^2 = x^2-4x+29

the equation is x^2-4x+29 = 0